show that surface a integral =48pi where a=4xi-2y^2j+z^2k and s is curved surface of cylinder x^2+y^2=4 bounded by plane z=0 to 3
1
Expert's answer
2020-05-07T15:48:36-0400

Let "x=2\cos\varphi" and "y=2\sin\varphi", so we have "s=\{(\varphi,z)|0\le\varphi\le 2\pi, 0\le z\le 3\}"

"(x'_{\varphi},y'_{\varphi},z'_{\varphi})=(-2sin\varphi,2\cos\varphi,0)" and "(x'_z,y'_z,z'_z)=(0,0,1)".

We have that normal to the cylinder is "\vec{N}=(x'_{\varphi},y'_{\varphi},z'_{\varphi})\times(x'_z,y'_z,z'_z)="

"=(-2sin\varphi,2\cos\varphi,0)\times(0,0,1)=(2\cos\varphi,2\sin\varphi,0)"

So "\iint\limits_s \vec{a}d\vec{s}=\iint\limits_s (\vec{a},\vec{N})d\varphi dz=\int\limits_0^3\int\limits_0^{2\pi}(\vec{a},\vec{N})d\varphi dz"

We have "\vec{a}=(4x,-2y^2,z^2)=(8\cos\varphi,-8\sin^2\varphi,z^2)", so "(\vec{a},\vec{N})=((8\cos\varphi,-8\sin^2\varphi,z^2),(2\cos\varphi,2\sin\varphi,0))="

"=16\cos^2\varphi-16\sin^3\varphi"

Since "\cos^2\varphi=\frac{1+\cos 2\varphi}{2}" and "\sin^3\varphi=\frac{3\sin\varphi-\sin 3\varphi}{4}", we have "(\vec{a},\vec{N})=8+8\cos 2\varphi-12\sin\varphi+4\sin 3\varphi"

Then "\iint\limits_s \vec{a}d\vec{s}=\int\limits_0^3\int\limits_0^{2\pi}(\vec{a},\vec{N})d\varphi dz="

"\int\limits_0^3\int\limits_0^{2\pi}(8+8\cos 2\varphi-12\sin\varphi+4\sin 3\varphi)d\varphi dz="

"=\int\limits_0^316\pi dz=48\pi"